# Signal Processing and Machine Learning Algorithms for Precise Timing with PICOSEC Micromegas Detectors

### v1

@@ Abstract
High particle rates in current and future experiments make pile-up phenomena a critical issue for extracting useful information. In this context, timing can be important as the 4th dimension parameter for triggering or event reconstruction. The PICOSEC-Micromegas detector has been shown to offer precise timing of the order of tens of ps. In this work, novel signal processing algorithms are being developed and evaluated to demonstrate the technology’s ability for online precise timing. We propose, an algorithm based on Artificial Neural Networks (ANN). This algorithm uses a model to train the ANN. The performance of the different algorithms is evaluated using experimental data, resulting in a timing resolution of 18.3 0.6 ps, comparable to the standard analysis based on the Constant Fraction Discrimination technique. Additionally, an alternative algorithm using the charge of the pulse exceeding a threshold as a parameter to correct for systematic effects is reported.
@@ 1 Introduction
The PICOSEC detector, shown in Fig.1(a), is based on the Micromegas technology [1](hereafter called PICOSEC-MM), and includes a Cherenkov radiator and a photocathode. This design allows to an incoming relativistic charged particle to create Cherenkov photons inside the radiator, which convert to prompt photoelectrons in the photocathode. Additionally, the drift region has been reduced to the order of , compared to a typical value of 3-5 mm to decrease the probability of direct ionization in the active volume, whilst also allowing for pre-amplification avalanche in the drift region due to the increased electric field.
The data used for this study were collected at the Saclay Laser-matter Interaction Center (IRAMIS/SLIC, CEA) with the LYDIL laser [3, 2]. The measurements were performed using an optimized PICOSEC-MM prototype with drift gap, operated at a high drift field and filled with (80:10:10) gas mixture. The electric field for these tests was set to 44 kV/cm for the drift and 21 kV/cm for the amplification detector’s volumes. The photocathode consisted of a thin film (18 nm) of CsI deposited on a 10 nm Aluminum layer, serving as the cathode. The laser beam was split, with one part directed at the PICOSEC-MM detector and the other illuminating a high-precision photodiode, with 3 ps timing resolution, serving as a timing reference. The light beam that reached the detector was properly attenuated to control the number of photons hitting the PICOSEC photocathode, allowing us to study the detector’s response to either a single-photoelectron or few, prompt photoelectrons [4]. The Signal Arrival Time (SAT) of the PICOSEC-MM was measured relative to the photodiode’s signal on an event-by-event basis (hereafter called reference signal). Previous studies have shown that this optimized PICOSEC-MM detector has the capability of timing the arrival of a single photon with an accuracy better than 50 ps [5].
@@
Two types of data sets were used: the SPE-set, comprising waveforms of single photoelectrons, and the EXP-set, composed of waveforms of multiple photoelectrons. All waveforms were digitized by a fast (20 GSamples/s), high-bandwidth oscilloscope (2.5 GHz) after passing through a CIVIDEC current-sensitive preamplifier. Section 2 first introduces an alternative timing method, based on the time-over-threshold technique. This algorithm uses the integral of the pulse above multiple thresholds (i.e. the equivalent of the charge) to correct for systematics. Building on the insights gained from these analytical signal processing approaches, the SPE-set is then used to emulate PICOSEC-MM signals with multiple photoelectrons, as described in Section 3. After evaluating the performance of this model in Section 4, it is used to train an Artificial Neural Network (ANN) for timing the arrival of PICOSEC-MM waveforms, as detailed in Section 5. In all methods discussed, the performance of the timing algorithms is assessed using the EXP-set waveforms.
It should be noted that the ANN-based timing technique performs optimally when the PICOSEC signal is relatively unaffected by time-walk. Due to the thin gap design, the prototype used in this study has a minimal time-walk effect, typically of only a few picoseconds [6]. However, in cases where the signals experience more significant time-walk dependencies, the technique is no longer a good approximation, potentially degrading the timing resolution.
@@ 2 Alternative Timing Technique with Constant Threshold
In our effort to propose an efficient timing technique, for scalable electronic devices with the PICOSEC-MM and reduce the cost of the experimental set-up, we evaluated the Constant Fraction method [7] on extracting the timing resolution of the PICOSEC-MM detector, by using the minimum information required. The goal of the developed signal processing algorithm is to retain the timing results achieved by the Constant Threshold method that takes into account the full waveform. In this approach, instead of using the time over threshold measurements, we use the integrals of the e-peak when exceeding certain thresholds. Hereafter such integrals will be referred to as Charge over Threshold and symbolized as . This technique, like the Time over Threshold, suffers from systematics, introduced by the strong dependence of SAT on the pulse height. We prove that we can extract the experimental information that can to correct this effect, achieving precise timing accuracy [8, 9].
@@
This technique requires thresholds adjusted to the amplitude range of the available experimental data. For the EXP-set, 4 thresholds were selected: 100 mV, 200 mV, 400 mV and 600 mV. The two crossing points of the threshold with the rising and falling edges of the waveform are used to integrate this region, where available, and calculate each value of the corresponding pulse. To make available the information of the e-peak amplitude, read-out electronics to digitize the waveform are required. However, on a large-scale experimental set-up, this will imply a significant number of channels, which can skyrocket the cost of the experimental set-up. Introducing pulse integrators such as the NINO-chip [10], or other new generation front-end electronics, in addition to the discriminator device, can achieve the time resolution of the Constant Fraction Discrimination Technique, at a significantly lower cost.
@@
The behaviour of mean SAT, can be evaluated as a function of e-peak charge above the applied threshold (100 mV), Fig.3(a), parameterising the time walk effect, using a power law plus a constant term fit. Correcting the SAT for time walk using this fit results is presented with black points on Fig.3(b). The algorithm that provides results with the best time resolution uses the lowest threshold for defining the SAT, and the parameterization as a function of of the highest available threshold for time walk correction, on an event-by-event basis [8]. This signal-processing algorithm minimizes the experimental information to a few numbers provided by the pulse integrators.
@@
The event-by-event time walk correction using the calibration curve of the higher available crossing threshold requires calibration measurements before any analysis ( Fig.4 ). The analysis of the EXP-set, using the technique, is shown in Fig.6(a), resulting in a timing resolution performance of the detector, as the one presented in Fig. 6(b), with black points. This is similar to the one achieved with the CFD Technique, which is expressed with red points in the same Figure and signifies the successful performance of the timing algorithm. After correction for systematic effects, SAT distribution is symmetric, with an RMS of 18.30.5 ps. The same results were achieved with the full offline analysis with CFD reported also in [8].
@@ 3 An Emulation Model for PICOSEC-MM Pulses
While analytical signal processing algorithms provide valuable benchmarks for timing extraction, their complexity can limit applicability in real-time environments. To address this, previous studies, such as [9, 11], have shown that ANN techniques can be used for online estimation of the arrival time of PICOSEC-MM waveforms with accuracy similar to the results of the full offline analysis. However, these studies were limited by the size of the available ANN training set and therefore relied on k-fold cross-validation methods for analysis [12]. In this work, a method for simulating large sets of waveforms is developed to train the ANN and provide more conclusive results.
The fully digitized waveforms of the SPE-set are used to emulate waveforms of the EXP-set, on the response of the PICOSEC-MM to a few pes. As it has been previously shown (e.g.[2, 3]), the e-peak charge distribution of the single-pe waveform is well described by the Gamma probability distribution function (pdf) of Eq.3.1:
@@
Operating the detector in such a high gain (i.e. 44 at the drift and over 21 at the amplification stage, generated by 525/275 V ), while creating the two data-sets, leads to a very small dependence of the pulses arrival time on their size, meaning that there is almost no time walk effect, as can be seen in Fig.8 . Using the full offline analysis algorithm, we can extract the timing performance of the two data sets as a function of the e-peak charge. Fig.9(a), shows the SAT as a function of the e-peak charge, confirming the small effect of the time-walk on the single photoelectron data-set, and it’s corresponding resolution, as in Fig.9(b). Respectively, Fig.10 represents the SAT of the EXP-set and its corresponding resolution, as a function of the e-peak charge.
@@ 3.1 Synchronizing and Summing Single Photo-electron Waveforms
As in [13], we assume that the EXP-set comprises PICOSEC-MM waveforms corresponding to photoelectrons, where follows a Poisson distribution with a mean value equal to . Then, the e-peak charge distribution of the EXP-set waveforms should follow the following pdf:
@@
A multiplicity, N of single-photoelectron (single-pe) signals is chosen according to a Poisson distribution with a mean value of 7.8. N waveforms are randomly selected among those of the SPE-set. Each selected waveform is then time-shifted, using a common set of time coordinates for all the SPE-set, where zero corresponds to the reference time. To accurately estimate the relative position of each single-pe signal, a degree polynomial interpolation is applied between the digitization points of the leading edge. This interpolation ensures that the pulses are synchronized, allowing for the summation of these signals into an equivalent multi-pe pulse. Synchronizing the reference time for all pulses by setting them to zero, creates a consistent clock for triggering and digitization. However, this synchronization removes the natural time jitter. To replicate this, we artificially introduce a Gaussian shift of 50 ps (with a standard deviation of 25 ps) to each summed pulse. This can be seen in an example of a simulated pulse in Fig.12(a). An example application of the algorithm, using N = 5 single-pe waveforms, is shown in Fig. 12(b).
Using the bootstrap technique [14] and the Poisson Distribution as a metric, we can generate random samples of single-pes to create simulated multi-pe pulses. However, this method has the disadvantage of introducing a correlation between the generated pulses since the final pulses share similar single-pe waveforms.
@@ 4 Validation of the Emulation Model
The data set of emulated multi-pes waveforms is expected to be identical to the response of the PICOSEC-MM detector on the laser beam (EXP-set). and share the same timing properties as the EXP-set. EXP-set data have been fully analyzed offline, following the algorithm described in [15]. After applying time walk corrections, parameterized as a function of the SAT to the e-peak charge, results in a timing resolution of 18.30.2 ps.
The full offline analysis of PICOSEC-MM waveforms, requires the calculation and subtraction of the baseline offset, while the baseline RMS is also calculated on an event-by-event basis [16]. For the SPE-set, the baseline RMS distribution is presented with black in Fig.13(b). Thus, during the summing process, the generated pulses present an increasing RMS baseline noise (red-colored distribution in Fig.13(b)), analogous with the square root of the number of photoelectrons.
@@ 4.2 Evaluating the Timing Properties of the Emulation Model
The evaluation of the timing properties of the model and their comparison with the EXP-set using the same offline analysis method, as performed in [8], will ensure the validity of the model.
It is essential, for the analysis of the emulated pulses, to quantify the correlation between e-peak size and the number of photoelectrons. For this, the distribution of e-peak charge is divided by the distribution of e-peak charge, weighed with the number of photoelectrons. The result is plotted as a function of e-peak charge in Fig.15(a), while the errors are calculated from the e-peak charge distribution weighted with the squared number of photoelectrons. The same results can be obtained for an analysis using the e-peak amplitude as a parameter (Fig. 15(b)). Each plot is parameterized using a degree polynomial.
@@
where the error of photodiode’s precision () is approximately 3 ps[4].
@@ 5.1 Validation of the Artificial Neural Network’s Performance
In the present study, the Neural Network implementation was based on a Feed-Forward NN architecture [17]. In this context, the ANN consists of an input layer, two hidden layers with 64 neurons each, and an output layer. This model has been found to give adequate results and can be visualized in Fig.17(a). For all nodes, ReLU activation function [18] was used apart from the output layer, an essential part of regression tasks. In parallel, the mean squared error was used as a loss function.
@@
Waveforms fed in the ANN were like the one seen in Fig.17(b). A 100 mV threshold trigger is chosen, giving a selection criterion to the waveforms and providing a timestamp. This timestamp is used to determine the starting time point which is 1.5 ns before the crossing point with the threshold. A shift of the starting time with a Gaussian with =500 ps, for the ANN to be unbiased of the selected starting point is applied. Subsequently, this will define the total time window of 3.2 ns (64-bit digitization with a sampling rate of 50 ps) given to the input layer of the ANN (red-colored points). Various input lengths were tested, and it was found that the optimum performance was achieved when the input layer contained mostly the leading edge. Moreover, during the training period, the reference time was provided to the network as the target value.

### v3

@@ Abstract
High particle rates in current and future experiments make pile-up phenomena a critical issue for extracting useful information. In this context, timing serves as a crucial fourth dimension for triggering and event reconstruction. The PICOSEC Micromegas detector has demonstrated precise timing capabilities on the order of tens of picoseconds. In this work, we develop and evaluate novel signal-processing algorithms to demonstrate the detector’s potential for online precise timing. In particular, we propose an algorithm based on Artificial Neural Networks (ANNs), trained using a physically motivated model. The performance of the different algorithms is assessed using experimental data recorded during laser beam tests at the IRAMIS Facility of CEA Saclay. A timing resolution of , corresponding to photoelectrons in the gas volume, is achieved—comparable to results obtained with the standard Constant Fraction Discrimination (CFD) technique. Additionally, we present an alternative algorithm that uses the integrated charge of the pulse exceeding a defined threshold as a parameter to correct for systematic effects.
@@ 1 Introduction
The PICOSEC detector, shown in Fig.1(a), is based on the Micromegas technology [GIOMATARIS](hereafter called PICOSEC-MM), and includes a Cherenkov radiator and a photocathode. This design allows to an incoming relativistic charged particle to create Cherenkov photons inside the radiator, which convert to prompt photoelectrons in the photocathode. Additionally, the drift region has been reduced to the order of , compared to a typical value of 3-5 mm to decrease the probability of direct ionization in the active volume, whilst also allowing for pre-amplification avalanche in the drift region due to the increased electric field.
The data used for this study were collected at the Saclay Laser–Matter Interaction Center (IRAMIS/SLIC, CEA) using the LYDIL laser system [BORTFELDT2021165049, BORTFELDT2018317]. The facility employs a mode-locked Ti:Sapphire laser (MIRA 900 555Coherent. Mira Optima 900 F Data Sheet (2002), pumped by VERDI V10666Coherent. Verdi Family Data Sheet (2011)) delivering pulses of (FWHM) duration at 800 nm with a 76 MHz repetition rate. The wavelength was subsequently converted to 265 nm through nonlinear frequency conversion using a barium borate (BBO) crystal, yielding ultraviolet pulses of approximately 150–200 fs (FWHM) at the photocathode. The measurements were performed using an optimized PICOSEC-MM prototype with drift gap, operated at a high drift field and filled with (80:10:10) gas mixture. The electric field for these tests was set to 44 kV/cm for the drift and 21 kV/cm for the amplification detector’s volumes. The photocathode consisted of a thin film (18 nm) of CsI deposited on a 10 nm Aluminum layer, serving as the cathode. The laser beam was split, with one part directed at the PICOSEC-MM detector and the other illuminating a high-precision photodiode, with 3 ps timing resolution, serving as a timing reference. The light beam that reached the detector was properly attenuated to control the number of photons hitting the PICOSEC photocathode, allowing us to study the detector’s response to either a single-photoelectron or few, prompt photoelectrons [sohl:tel-03167728]. The Signal Arrival Time (SAT) of the PICOSEC-MM was measured relative to the photodiode’s signal on an event-by-event basis (hereafter called reference signal). Previous studies have shown that this optimized PICOSEC-MM detector has the capability of timing the arrival of a single photon with an accuracy better than 50 ps [Sohl_2020].
@@
Two types of data sets were used: the SPE-set, comprising waveforms of single photoelectrons ( approximately 16k waveforms dataset), and the EXP-set (approximately 12k waveforms dataset), composed of waveforms of multiple photoelectrons. All waveforms were digitized by a fast (20 GSamples/s), high-bandwidth oscilloscope (2.5 GHz) after passing through a CIVIDEC current-sensitive preamplifier. Section 2 first introduces an alternative timing method, based on the time-over-threshold technique. This algorithm uses the integral of the pulse above multiple thresholds (i.e. the equivalent of the charge) to correct for systematics. Building on the insights gained from these analytical signal processing approaches, the SPE-set is then used to emulate PICOSEC-MM signals with multiple photoelectrons, as described in Section 3. After evaluating the performance of this model in Section 4, it is used to train an Artificial Neural Network (ANN) for timing the arrival of PICOSEC-MM waveforms, as detailed in Section 5. In all methods discussed, the performance of the timing algorithms is assessed using the EXP-set waveforms.
It should be noted that the ANN-based timing technique performs optimally when the PICOSEC signal is relatively unaffected by time-walk. Due to the thin gap design, the prototype used in this study has a minimal time-walk effect, typically of only a few picoseconds [utrobicic2024singlechannelpicosecmicromegas]. However, in cases where the signals experience more significant time-walk dependencies, the technique is no longer a good approximation, potentially degrading the timing resolution.
@@ 2 Alternative Timing Technique with Constant Threshold
In our effort to propose an efficient timing technique, for scalable electronic devices with the PICOSEC-MM and reduce the cost of the experimental set-up, we evaluated the Constant Fraction method [georgakopoulou2018100] on extracting the timing resolution of the PICOSEC-MM detector, by using the minimum information required. The goal of the developed signal processing algorithm is to retain the timing results achieved by the Constant Threshold method that takes into account the full waveform. In this approach, instead of using the time over threshold measurements, we use the integrals of the e-peak when exceeding certain thresholds. Hereafter such integrals will be referred to as Charge over Threshold and symbolized as . This technique, like the Time over Threshold, suffers from systematics, introduced by the strong dependence of SAT on the pulse height. We prove that we can extract the experimental information that can to correct this effect, achieving precise timing accuracy [kallitsopoulou2021development, Manthos_2021].
@@
This technique requires thresholds adjusted to the amplitude range of the available experimental data. For the EXP-set, 4 thresholds were selected: 100 mV, 200 mV, 400 mV and 600 mV. The two crossing points of the threshold with the rising and falling edges of the waveform are used to integrate this region, where available, and calculate each value of the corresponding pulse. To make available the information of the e-peak amplitude, read-out electronics to digitize the waveform are required. However, on a large-scale experimental set-up, this will imply a significant number of channels, which can skyrocket the cost of the experimental set-up. Introducing pulse integrators such as the NINO-chip [NINO], or other new generation front-end electronics, in addition to the discriminator device, can achieve the time resolution of the Constant Fraction Discrimination Technique, at a significantly lower cost.
@@
The behavior of the mean SAT can be evaluated as a function of e-peak charge above the applied threshold (100 mV), Fig. 3(a), parameterizing the time-walk effect using a power-law plus constant-term fit. Correcting the SAT for time walk using this fit is shown by the black points in Fig. 3(b). The algorithm that provides the best time resolution uses the lowest threshold to define the SAT, and the parameterization as a function of of the highest available threshold for the time-walk correction, on an event-by-event basis [kallitsopoulou2021development]. Each pulse is corrected using the parameterization corresponding to the highest threshold crossed by that specific signal. Pulses that cross only one or two thresholds are still included, and no events are discarded from the dataset. This approach ensures that every waveform contributes to the final statistics while exploiting the most accurate correlation available for its amplitude range. This signal-processing algorithm minimizes the experimental information to a few numbers provided by the pulse integrators.
@@
The event-by-event time walk correction using the calibration curve of the higher available crossing threshold requires calibration measurements before any analysis ( Fig.4 ). The analysis of the EXP-set, using the technique, is shown in Fig.6(a), resulting in a timing resolution performance of the detector, as the one presented in Fig. 6(b), with black points. This is similar to the one achieved with the CFD Technique, which is expressed with red points in the same Figure and signifies the successful performance of the timing algorithm. After correction for systematic effects, SAT distribution is symmetric, with an RMS of 18.30.5 ps. The same results were achieved with the full offline analysis with CFD reported also in [kallitsopoulou2021development].
@@ 3 An Emulation Model for PICOSEC-MM Pulses
3.1 Characterization of Single-Photoelectron Waveforms Properties
While analytical signal processing algorithms provide valuable benchmarks for timing extraction, their complexity can limit applicability in real-time environments. To address this, previous studies, such as [Manthos_2021, Tsiamis_2022], have shown that ANN techniques can be used for online estimation of the arrival time of PICOSEC-MM waveforms with accuracy similar to the results of the full offline analysis. However, these studies were limited by the size of the available ANN training set and therefore relied on k-fold cross-validation methods for analysis [francois]. In this work, a method for simulating large sets of waveforms is developed to train the ANN and provide more conclusive results.
The fully digitized waveforms of the SPE-set are used to emulate waveforms of the EXP-set, on the response of the PICOSEC-MM to a fewphotoelectrons. As it has been previously shown (e.g.[BORTFELDT2018317, BORTFELDT2021165049]), the e-peak charge distribution of the single-pe waveform is well described by the Gamma probability distribution function (pdf) of Eq.3.1:
@@ 3.1 Characterization of Single-Photoelectron Waveforms Properties
Operating the detector in such a high gain (i.e. 44 at the drift and over 21 at the amplification stage, generated by 525/275 V ), while creating the two data-sets, leads to a very small dependence of the pulses arrival time on their size, meaning that there is almost no time walk effect, as can be seen in Fig.8. Although the voltage settings used for the data presented in this work are not included here, the figure illustrates the general trend that higher electric fields lead to a reduced time-walk effect. Using the full offline analysis algorithm, we can extract the timing performance of the two data sets as a function of the e-peak charge. Fig.9(a), shows the SAT as a function of the e-peak charge, confirming the small effect of the time-walk on the single photoelectron data-set, and it’s corresponding resolution, as in Fig.9(b). Respectively, Fig.10 represents the SAT of the EXP-set and its corresponding resolution, as a function of the e-peak charge.
@@ 3.2 Synchronizing and Summing Single Photo-electron Waveforms
As in [Manthos_2020], we assume that the EXP-set comprises PICOSEC-MM waveforms corresponding to photoelectrons, where follows a Poisson distribution with a mean value equal to . Then, the e-peak charge distribution of the EXP-set waveforms should follow the following pdf:
@@
A multiplicity, N of single-photoelectron (single-pe) signals is chosen according to a Poisson distribution with a mean value of 7.8. N waveforms are randomly selected among those of the SPE-set. Each selected SPE waveform is time-aligned to a common reference (zero) using a 3rd-degree polynomial interpolation of the leading edge, and the aligned waveforms are summed to form an equivalent multi-photoelectron pulse. This exact alignment, however, artificially removes the natural timing spread that exists between pulses and the finite sampling-phase/trigger jitter of the digitizer: with perfect alignment, every summed pulse would have identical digit values at each sampling point. To restore realistic conditions, we introduce a stochastic time-shift: each summed pulse is shifted by a random Gaussian offset with standard deviation (corresponding roughly to range), thereby emulating the combined effects of photoelectron arrival-time spread and digitizer/trigger sampling jitter. An example before and after applying this jitter is shown in Fig. 12(b).
Using the bootstrap technique [ramachandran2009mathematical] and the Poisson Distribution as a metric, we can generate random samples of single-pes to create simulated multi-pe pulses (approximately 100k events data set). However, this method has the disadvantage of introducing a correlation between the generated pulses since the final pulses share similar single-pe waveforms.
@@ 4 Validation of the Emulation Model
The data set of emulated multi-pes waveforms is expected to be identical to the response of the PICOSEC-MM detector on the laser beam (EXP-set). and share the same timing properties as the EXP-set. EXP-set data have been fully analyzed offline, following the algorithm described in [Aune:2020mwh]. After applying time walk corrections, parameterized as a function of the SAT to the e-peak charge, results in a timing resolution of 18.30.2 ps, shown in Fig.13(a).
The full offline analysis of PICOSEC-MM waveforms requires the calculation and subtraction of the baseline offset, while the baseline RMS is also determined on an event-by-event basis [Bortfeldt2019]. For each waveform, the baseline is estimated within a signal-free time window of at least 80 ns preceding the pulse. The amplitude distribution in this region is fitted with a Gaussian function, whose mean value defines the baseline offset and whose standard deviation corresponds to the baseline RMS. This procedure corrects baseline drifts and accounts for fluctuations in the electronic noise level on an event-by-event basis. For the SPE-set, the baseline RMS distribution is presented with black in Fig.13(b). Thus, during the summing process, the generated pulses present an increasing RMS baseline noise (red-colored distribution in Fig.13(b)), analogous with the square root of the number of photoelectrons.
@@ 4.2 Evaluating the Timing Properties of the Emulation Model
The evaluation of the timing properties of the model and their comparison with the EXP-set using the same offline analysis method, as performed in [kallitsopoulou2021development], will ensure the validity of the model.
It is essential, for the analysis of the emulated pulses, to quantify the correlation between the e-peak size and the corresponding number of photoelectrons. To determine this relation, two distributions are constructed from the simulated dataset: (i) the total e-peak charge distribution, and (ii) the same distribution where each event is weighted by its number of photoelectrons, . The ratio of the weighted to the unweighted distributions provides the mean number of photoelectrons as a function of the e-peak charge, as shown in Fig. 15(a). The statistical uncertainties are derived from the weighted charge distribution and scaled with . A similar analysis can be performed using the e-peak amplitude instead of the charge (Fig. 15(b)). Each resulting dependence is parameterized with a fourth-degree polynomial for subsequent use in the waveform reconstruction.
@@
where the error of photodiode’s precision () is approximately 3 ps[sohl:tel-03167728].
@@ 5.1 Validation of the Artificial Neural Network’s Performance
In the present study, the Neural Network implementation was based on a Feed-Forward NN architecture [goodfellow2016deep]. In this context, the ANN consists of an input layer, two hidden layers with 64 neurons each, and an output layer. This model has been found to give adequate results and can be visualized in Fig.17(a). For all nodes, ReLU activation function [agarap2019deep] was used apart from the output layer, an essential part of regression tasks. In parallel, the mean squared error was used as a loss function.
@@
The predictive capabilities of the ANN were first studied with experimental data using the k-fold cross-validation technique, which was already mentioned in the introduction of Section 3 [francois].
@@
Waveforms fed in the ANN were like the one seen in Fig.17(b). Red points denote the digitized information presented to the ANN, in this case the 64 digits around the trigger point-corresponding to 3.2 ns time interval. A 100 mV threshold trigger is chosen, giving a selection criterion to the waveforms and providing a timestamp. This timestamp is used to determine the starting time point which is 1.5 ns before the crossing point with the threshold. A shift of the starting time with a Gaussian with =500 ps, for the ANN to be unbiased of the selected starting point is applied. Subsequently, this will define the total time window of 3.2 ns (64-bit digitization with a sampling rate of 50 ps) given to the input layer of the ANN (red-colored points). Various input lengths were tested, and it was found that the optimum performance was achieved when the input layer contained mostly the leading edge. Moreover, during the training period, the reference time was provided to the network as the target value.